Which of the following is a better buy? A 16 inch pizza for $12 or two 6 inch pizzas for $9
Answer by mananth(16949)

the one which costs less per in^2 is better buy
..
16 inch pizza
area = pi*8*8
area = 201 in^2
...
201 in^2 pizza costs $12
1 in^2 = 12/201
1 in^2 = 5.97 = 6 cents
...
6" dia pizza
area = pi*3*3
area = 9pi
area = 28.27 in^2
2 pizzas = 2*28.27
= 56.54 in^2
...
56.54 in^2 ----- 9
1 in^2= 9/56.54
=0.159
=0.16
= 16 cents